OpenStudy (anonymous):

Can someone check my answer. Solve the equation. 6.9x - 8.4 = 4.02 I got 1.38

OpenStudy (anonymous):

no sorry thats wrong

jigglypuff314 (jigglypuff314):

6.9x - 8.4 = 4.02 add 8.4 to both sides +8.4 +8.4 6.9x = 12.42 divide both sides by 6.9 /6.9 /6.9 x = 12.42 divided by 6.9 =

OpenStudy (anonymous):

thanks @jigglypuff314 and @Aramaiel for helping